OpenStudy (anonymous):

Write 26,340,000 in scientific notation.....these scientific notations r killing me

Parth (parthkohli):

Shift decimals!

OpenStudy (anonymous):

ok how

Parth (parthkohli):

\[26340000 \implies 26340000 \times 1 \implies 2.634000.0 \times 10 \implies 263400.00 \times 100 \]\[\implies 26340.000 \times 1000\implies 2634.0000 \times 10,000 \implies 263.4 \times 100,000\]\[\implies 26.34 \times 1,000,000 \implies 2.634 \times 10,000,000 \]